CAVITY RINGDOWN SPECTROSCOPY AND KINETICS OF BUTOXY ISOMERIZATION: DETECTION OF THE Ã-X̃ BAND OF HOC_{4}H_{8}OO

\underline{MATTHEW K. SPRAGUE}^{a}, MITCHIO OKUMURA, California Institute of Technology, Division of Chemistry, MC 127-72, Pasadena, CA 91125; and STANLEY P. SANDER, Jet Propulsion Laboratory, California Institute of Technology, MS 183-901, Pasadena, CA 91109.

Alkoxy radicals are atmospherically important species, playing a direct role in the HO_{x} and NO_{x} cycles that affect tropospheric air pollution. Alkoxy radicals that can form a six membered transition state can isomerize into a hydroxyalkyl radical. In the presence of O_{2}, a rapid association reaction occurs, forming a hydroxyalkylperoxy radical. The n-butoxy radical is the smallest alkoxy to undergo isomerization, forming the γ-hydroxybutylperoxy radical (HOC_{4}H_{8}OO). Direct detection of HOC_{4}H_{8}OO opens the door to measuring alkoxy reaction kinetics with higher precision than previous end product studies.^{b} In this talk, we report the first detection of the Ã-X̃ electronic spectrum of the HOC_{4}H_{8}OO radical. The spectrum is similar in shape to the Ã-X̃ spectrum of n-butyl peroxy:^{c} a broad spectroscopic band due to the multiple molecular conformers that are present. We also use the Ã-X̃ band to measure the alkoxy isomerization kinetics relative to the alkoxy radical’s reaction with O_{2}, similar to the previous kinetics studies.

aSupport from the NDSEG Fellowship, California Air Resources Board Contracts 03-333 and 07-730, and NASA Upper Atmosphere Research Program Grants NAG5-11657, NNG06GD88G and NNX09AE21G are gratefully acknowledged

bR. Atkinson, D. L. Baulch, R. A. Cox, J. N. Crowley, R. F. Hampson, R. G. Hynes, M. E. Jenkin, M. J. Rossi, and J. Troe J.

Atmos. Chem. and Phys. 2006, 6, 3625.

cB. G. Glover and T. A. Miller. J. Phys. Chem. A 2005, 109, 11191.